OpenStudy (anonymous):

MEDAL GIVEN TO CORRECT ANSWER. A soccer player runs across a soccer field. a. If the soccer player runs with a speed of 4.6 m/s, how long does it take him to run 60 m? b. It takes the soccer player 3 seconds to reduce his speed to 3.1 m/s. What is his acceleration? c. The soccer player is running at a speed of 4 m/s when he slides on the ground, coming to a stop in a distance of 2 m. What was his acceleration? d. The soccer player accelerates from rest at a rate of 3.1 m/s2. How far does he run in 1.5 seconds?

OpenStudy (ybarrap):

Here are the equations you need for each question: a) $$ s=vt\\ \implies t=\cfrac{s}{t} $$ b) $$ a=\cfrac{v_f-v_i}{\Delta t} $$ c) $$ s=\cfrac{v_i+v_f}{2}\Delta t\\ =\cfrac{v_i+0}{2}\Delta t\\ \implies \Delta t=\cfrac{s}{\cfrac{v_i}{2}}\\ a=\cfrac{v_f-v_i}{\Delta t}\\ a=\cfrac{0-v_i}{\Delta t} $$ d) $$ s=\cfrac{v_i+v_f}{2}\Delta t\\ =\cfrac{0+v_f}{2}\Delta t $$ Now just fill in the variables with your givens. Make sense?

OpenStudy (anonymous):

you just confused me even more, tbh.

OpenStudy (anonymous):

Time is distance divided by speed. Acceleration is change in velocity divided by time. Stopping a velocity v requires a negative acceleration such that v squared = twice the distance times the acceleration distance from stopped position is (1/2)(acceleration)(ime)^2

OpenStudy (ybarrap):

Plugging in values, we get: Question a $$ s=vt\\ \implies t=\cfrac{s}{v}\\ t=\cfrac{60}{4.6} \text{s} $$ Question b $$ a=\cfrac{v_f-v_i}{\Delta t}\\ =\cfrac{3.1-4.6}{3}\cfrac{\text{m }}{\text{s}^2} $$ Question c $$ s=\cfrac{v_i+v_f}{2}\Delta t\\ =\cfrac{v_i+0}{2}\Delta t\\ \implies \Delta t=\cfrac{s}{\cfrac{v_i}{2}}\\ \Delta t=\cfrac{2}{\cfrac{3}{2}}=\cfrac{1}{3}\text{ s}\\ a=\cfrac{v_f-v_i}{\Delta t}\\ a=\cfrac{0-v_i}{\Delta t}\\ a=\cfrac{0-4}{\cfrac{1}{3}}=-12\cfrac{\text{m }}{\text{s}^2} $$ Question d $$ s=\cfrac{v_i+v_f}{2}\Delta t\\ =\cfrac{0+v_f}{2}\Delta t\\ =\cfrac{3.1}{2}1.5\text{ m}\\ $$